OpenStudy (anonymous):

find integral of ysqrt(y+3)dy

OpenStudy (anonymous):

\[\int\limits_{}^{} y \sqrt{y+3}dy\]

OpenStudy (anonymous):

put u=sqrt(y+3)

OpenStudy (anonymous):

ummm which method are you using?

OpenStudy (anonymous):

substitution if i remember correctly,should work here

OpenStudy (anonymous):

ummm but the derivative of the outside is not y. Err, I mean 1 is there technically, but then what happened to the y up front?

OpenStudy (anonymous):

write it in terms of u,it;ll be y=u^2-3

OpenStudy (anonymous):

wait I'm sorry, do you mean u^2/3?

OpenStudy (anonymous):